OpenStudy (anonymous):

peter moran bought a dell laptop computer priced at $699. he put down 30 %. the amount of down payment is what

OpenStudy (mathstudent55):

Since the amount he put down is 30% of the price, he put down 30% of $699. 30% of $699 means \(30\% \times $699 \) % = \(\dfrac{1}{100} = 0.01\) So, \(30\% \times $699 \) \( = 30 \times 0.01 \times $699 \)

OpenStudy (anonymous):

209.70 thank you

OpenStudy (mathstudent55):

Correct. You're welcome.